OpenStudy (anonymous):

Can someone explain how to do this problem? Calculate the wavelength in nm that would be emitted by an electron in a helium ion (He+) falling from n = 6 to n = 4 Thanks

OpenStudy (cuanchi):

it is a two steps calculation, first you have to calculate the energy released by the electron falling from 6 to 4 and then calculate the wavelength associated to that energy

OpenStudy (cuanchi):

\[\Delta E= -2.18 \times10^{-18}J(\frac{ 1 }{ n _{f}^{2} }-\frac{ 1 }{ n _{i}^{2} })\] \[E= h c/\lambda \] \[\lambda =E/(hc)\]